# Conformal Prediction Sets Can Cause Disparate Impact

### v1

@@ Abstract
Although conformal prediction is a promising method for quantifying the uncertainty of machine learning models, the prediction sets it outputs are not inherently actionable. Many applications require a single output to act on, not several. To overcome this, prediction sets can be provided to a human who then makes an informed decision. In any such system it is crucial to ensure the fairness of outcomes across protected groups, and researchers have proposed that Equalized Coverage be used as the standard for fairness. By conducting experiments with human participants, we demonstrate that providing prediction sets can increase the unfairness of their decisions. Disquietingly, we find that providing sets that satisfy Equalized Coverage actually increases unfairness compared to marginal coverage. Instead of equalizing coverage, we propose to equalize set sizes across groups which empirically leads to more fair outcomes.
@@ 1 Introduction
Figure 1: We measure the increase in accuracy per-group when prediction sets are provided to humans compared to the control population without model assistance. Disparate impact is the maximum difference of increases between groups, and should be minimized for fairness. Prediction sets do not benefit all groups equally, while sets with Equalized Coverage (Conditional) lead to the most unfair outcomes. Statistical analyses and significance are presented in Section 6.
@@
Still, there are many cases where decisions are not left to a model alone, but require a human-in-the-loop. For example, society does not yet use machine learning models to make most medical diagnoses – we prefer human doctors for reasons of safety, accountability, and trust. Nevertheless, there are ample opportunities for people like doctors to use machine learning tools to improve their decisions. As uncertainty quantification is a crucial component for trust, it has been argued that CP sets are a natural fit for such applications (Lu et al., 2022). Indeed, research has shown that providing conformal sets to humans increases their accuracy on decision tasks (Straitouri & Gomez Rodriguez, 2024; Zhang et al., 2024; Cresswell et al., 2024; De Toni et al., 2024).
@@ 2.1 Conformal set predictors
CP (Vovk et al., 2005; Shafer & Vovk, 2008) is one of the leading approaches to uncertainty quantification in machine learning (Soize, 2017; Abdar et al., 2021), and is aligned with efforts to make models more trustworthy. Other uncertainty quantification approaches often rely on strong assumptions which are unlikely to hold in practice (Gal & Ghahramani, 2016; Lakshminarayanan et al., 2017), or require modifications to the model architecture (Neal, 2012). In contrast, CP applies to black-box models, has no dependence on training data, is distribution free, is valid in finite samples, and assumes only exchangeability between calibration and test data – a weaker assumption than being i.i.d. which makes it realistic in practice (Vovk et al., 1999; Angelopoulos & Bates, 2021).
@@
CP achieves coverage by allowing the set size $|\mathcal{C}(x)|$ to vary depending on the model’s heuristic confidence, e.g., the softmax outputs of $f$. Larger sets indicate greater model uncertainty about $x$. To move from $f$ to $\mathcal{C}$, CP carries out a calibration step on a held-out dataset $\mathcal{D}_{\text{cal}}$ of $n_{\text{cal}}$ datapoints drawn independently from $\mathbb{P}$. First, one defines a conformal score function $s:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ with larger scores indicating worse agreement between $x$ and $y$ according to the heuristic uncertainty notion, and then computes $s$ on all datapoints $(x,y)\in\mathcal{D}_{\text{cal}}$. After sorting the scores, one computes the $\tfrac{\lceil{(n+1)(1-\alpha)}\rceil}{n}$ quantile which we will denote $\hat{q}$. With the calibration step done, CP sets can be generated for a test datapoint $x_{\text{test}}$ as
$\mathcal{C}_{\hat{q}}(x_{\text{test}})=\{y\mid s(x_{\text{test}},y)<\hat{q}\}.$ (2)
For any classifier $f$, and any score function $s$, the CP sets $\mathcal{C}_{\hat{q}}$ will obey the coverage guarantee (Equation 1) under the exchangeability assumption.
@@ 2.3 Non-conformal set predictors
$\mathcal{C}_{q_{k}}(x_{\text{test}})=\{y\mid f(x_{\text{test}})_{y}>q_{k}\}.$ (4)
@@ 2.4 Fairness of set predictors
$\delta_{t,a}=\mathrm{acc}_{t}(x\in\mathcal{D}\mid g=a)-\mathrm{acc}_{\mathrm{control}}(x\in\mathcal{D}\mid g=a),$ (6)
@@
$\Delta_{t}=\max_{a,b\in\mathcal{G}}(\delta_{t,a}-\delta_{t,b}),$ (7)
@@
$\Delta_{\text{Cov}}=\max_{a,b\in\mathcal{G}}\left(\text{Cov}(x\in\mathcal{D}\mid g=a)-\text{Cov}(x\in\mathcal{D}\mid g=b)\right).$ (8)
@@ 4.1 Hypotheses
On its own this tradeoff is not directly connected to the substantive fairness of using prediction sets. However, it has been observed in previous studies that it is set size, not coverage, which has the greatest influence on human accuracy, and therefore on outcomes (Cresswell et al., 2024). Together, these facts imply that equalizing coverage will create more imbalanced set sizes which will propagate into more biased human performance. The proposed mechanism by which equalizing coverage with Mondrian (conditional) CP causes unfairness is illustrated in Figure 2.
@@ 4.2 Method
Alongside empirical estimates of the standard errors, it is straightforward to derive odds ratios (ORs) from the GEE model. ORs allow us to estimate how much more likely a participant in treatment $t$ is to give the correct answer than if they were in the control, which is another way of expressing the expected accuracy improvement due to $t$ (c.f. Equation 6). For the sake of clarity, we give the formal description below. Let $p_{t,a}=\hat{\mathbb{P}}(\mathrm{correct}=1\mid\mathrm{treat}=t,\mathrm{group}=a)$ be the probability of a correct answer under the GEE model, conditional on treatment and group. For treatment $t$, and for every protected group $a$, the OR of $t$ versus control is given by
$\mathbf{OR}_{t,a}=\frac{p_{t,a}/(1-p_{t,a})}{p_{\mathrm{control},a}/(1-p_{\mathrm{control},a})}.$ (10)
@@
To assess the disparate impact caused by any treatment for Hypothesis 1, we can compare ORs between groups. We define the ratio of ORs (RORs) for treatment $t$ and groups $a$ and $b$ as $\mathbf{ROR}_{t,a,b}=\mathbf{OR}_{t,a}/\mathbf{OR}_{t,b}$. When $\mathbf{ROR}_{t,a,b}\approx 1$, treatment $t$ provides the same advantage over the control for group $a$ as it does for group $b$ which is a fair outcome. Hence, a value of $\mathbf{ROR}_{t,a,b}$ greater than 1 for some pair $(a,b)$ indicates that $t$ causes disparate impact, and so we report the maximum ROR between pairs of groups as a descriptive statistic (c.f. Equation 7),
@@ 5.1 Tasks, Datasets, and Models
Additional details on tasks, datasets, models, and set prediction methods are given in Appendix A. Our code is provided as supplemental material.
@@ 5.2 Experiment Design
During each experiment, participants were first shown a consent form detailing how their data would be collected and used. Then, participants were introduced to the task and trained on 20 practice trials which were not used in our analysis. The testing phase proceeded with $M=50$ more trials, an example of which is shown in Figure 3 for FACET, while detailed task descriptions can be found in Appendix B. For each trial, participants were presented with a datapoint $x$ along with all $m$ class labels and were asked to classify $x$. For the avg-$k$, marginal, and conditional treatments participants were also shown a prediction set, accompanied by the expected coverage. There was no time limit for responses, however for FACET the stimulus was only displayed for one second to increase task difficulty. After the class was selected for each trial, the correct answer was revealed.
@@ 6.1 Disparate Impact Measurement In Human Subject Experiments
Hypothesis 1 supposes that biases from a model can propagate to humans using prediction sets as an aid and result in disparate impact. Table 1 demonstrates that the models we used have some biases across groups, with non-zero $\Delta_{\text{Top-1}}$ values, which translated to non-zero $\Delta_{\text{Cov}}$ for the avg-$k$ and marginal sets. Conditional sets do equalize coverage ($\Delta_{\text{Cov}}\approx 0$), but at the cost of greater discrepancies in set size, $\Delta_{\text{Size}}$. From Table 2 we see that in most cases prediction sets were useful to participants versus the control ($\mathbf{OR}_{t,a}>1$), supporting previous findings (e.g. Cresswell et al. (2024)). However, it is clear that the beneficial effect is not experienced equally across groups. For FACET, the Younger group had the highest model accuracy (Figure 5), and experienced the most improvement across all treatments. The Older and Unknown groups having lower model accuracy and minority representation (see Table 6 in Appendix A) saw less improvement, and even some harm with conditional sets where human performance decreased. Hence, all three treatments caused disparate impact ($\mathbf{maxROR}_{t}>1)$. Similar results are seen for BiosBias. Even though the model itself was much less biased (Table 1) and both Female and Male groups benefited when sets were provided, the amount of benefit was not equally shared. Both FACET and BiosBias show that providing prediction sets can cause disparate impact, in support of Hypothesis 1. RAVDESS provides an example that disparate impact need not always occur. Again, both groups benefited from each treatment, but for the avg-$k$ and marginal treatments the increase actually was commensurate across groups. For these two treatments the prediction sets did not have equal coverage, but did have close to the same average set sizes across groups.
@@ 6.2 Insights
By contrast, we see strong correlation between set size differences and $\Delta_{t}$. The most fair outcomes occur when set sizes are similar across all groups, as for RAVDESS-Marginal, or FACET-Avg-K. The conditional treatment which increases set size differences in its effort to equalize coverage has an undesirable effect on fairness. Similar to set size, there is also clear correlation between singleton frequency and $\Delta_{t}$. Cresswell et al. (2024) observed that human accuracy on tasks was highest when the model expressed certainty through singleton sets. Hence, groups with more singletons could expect higher accuracy, so it is plausible that equalizing how often singletons occur would encourage fair outcomes, which is what we observe.
In summary, our results show that simply equalizing coverage, as is the aim of the conditional treatment, is not an effective way to promote fairness. Instead, focusing on Equalized Set Size or Equalized Singleton Frequency could be more useful standards of fairness for set predictors.
@@ Appendix A Implementation Details
Our code for curating datasets and performing set prediction is available at github.com/layer6ai-labs/conformal-prediction-fairness. In this section we detail the steps that were taken.
@@ A.1 Dataset Pre-processing
For RAVDESS (Livingstone & Russo, 2018), a dataset of audio-visual recordings where actors express emotions, we worked with 1,440 audio-only samples exhibiting $m=8$ emotions: [Happy, Angry, Calm, Fearful, Neutral, Disgust, Sad and Surprised]. The dataset contains recordings from 24 actors, 12 female and 12 male, so we used emotion as the class label and gender as the protected group. Each actor recorded 60 snippets of approximately 4 seconds in length each. The 60 recordings are broken down as 2 re-recordings each of 2 intensities each of 2 statements of 8 emotions. However, the “Neutral” emotion used only a single intensity. We split the dataset into 240 for calibration-validation, 840 samples for calibration, and 360 for testing using stratified sampling over both emotion and gender. Given that each split was treated the same way they can be considered i.i.d. for conformal predictions. To process the audio samples, we used a wav2Vec2 model Baevski et al. (2020) available on Huggingface (Fadel, 2023). We resampled each file to 16kHz as the base model was pre-trained on that frequency. For audio with multiple channels, we changed it to mono by averaging channels. The model’s feature extractor then turned the raw audio signals into feature vectors ready for the model to use, which were subsequently padded to ensure consistent input length for batch processing by the model. RAVDESS is released under a Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International license.
@@ Appendix B Human Subject Experiments
Ethics Considerations Throughout our research, we took steps to ensure that research ethics were upheld. We acknowledge that ethical considerations are extremely important for any research involving human subjects, and followed the ethical standards of our institution. Although it does not have an internal review board (IRB) process, the steps we took to meet its standards included: understanding the protocols and procedures for ethical research at our institution; curating datasets that only contained content suitable for showing to participants; performing the experiments on ourselves as authors; providing participants with a clear consent form that detailed the data that would be collected as well as how it would be used and how long it would be retained; paying all recruits a fair wage; and ensuring we were aware of potential biases in the data we collected by reviewing the demographic distribution of recruits (Table 8).

### v2

@@ Abstract
Conformal prediction is a statistically rigorous method for quantifying uncertainty in models by having them output sets of predictions, with larger sets indicating more uncertainty. However, prediction sets are not inherently actionable; many applications require a single output to act on, not several. To overcome this limitation, prediction sets can be provided to a human who then makes an informed decision. In any such system it is crucial to ensure the fairness of outcomes across protected groups, and researchers have proposed that Equalized Coverage be used as the standard for fairness. By conducting experiments with human participants, we demonstrate that providing prediction sets can lead to disparate impact in decisions. Disquietingly, we find that providing sets that satisfy Equalized Coverage actually increases disparate impact compared to marginal coverage. Instead of equalizing coverage, we propose to equalize set sizes across groups which empirically leads to lower disparate impact.
@@ 1 Introduction
Figure 1: We measure the increase in accuracy per-group from using prediction sets compared to the control population without model assistance. Disparate impact is the maximum difference of increases between groups, and should be minimized for fairness (Equation 7). Prediction sets do not benefit all groups equally, while sets with Equalized Coverage (Conditional) lead to the most unfair outcomes. Statistical analyses and significance are presented in Section 6.
@@
Still, there are many cases where decisions are not left to a model alone, but require a human in the loop. For example, society does not yet use machine learning models to make most medical diagnoses – we prefer human doctors for reasons of safety, accountability, and trust. Nevertheless, there are ample opportunities for people like doctors to use machine learning tools to improve their decisions. As uncertainty quantification is a crucial component for trust, it has been argued that CP sets are a natural fit for such applications (Lu et al., 2022). Indeed, research has shown that providing conformal sets to humans increases their accuracy on decision tasks (Straitouri & Gomez Rodriguez, 2024; Zhang et al., 2024; Cresswell et al., 2024; De Toni et al., 2024).
@@ 2.1 Conformal set predictors
CP (Vovk et al., 2005; Shafer & Vovk, 2008) is one of the leading approaches to uncertainty quantification in machine learning (Soize, 2017; Abdar et al., 2021), and is aligned with efforts to make models more trustworthy. Other uncertainty quantification approaches often rely on strong assumptions which are unlikely to hold in practice (Gal & Ghahramani, 2016; Lakshminarayanan et al., 2017), or require modifications to the model architecture (Neal, 2012). In contrast, CP applies to black-box models, has no dependence on training data, is distribution free, is valid in finite samples, and assumes only that test data is drawn from the same distribution as calibration data (Vovk et al., 1999; Angelopoulos & Bates, 2021).
@@
CP achieves coverage by allowing the set size $|\mathcal{C}(x)|$ to vary depending on the model’s heuristic confidence, e.g., the softmax outputs of $f$. Larger sets indicate greater model uncertainty about $x$. To move from $f$ to $\mathcal{C}$, CP carries out a calibration step on a held-out dataset $\mathcal{D}_{\text{cal}}$ of $n_{\text{cal}}$ datapoints drawn independently from $\mathbb{P}$. First, one defines a conformal score function $s:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ with larger scores indicating worse agreement between $x$ and $y$ according to the heuristic uncertainty notion, and then computes $s$ on all datapoints $(x,y)\in\mathcal{D}_{\text{cal}}$. After sorting the scores, one computes the $\tfrac{\lceil{(n+1)(1-\alpha)}\rceil}{n}$ quantile which we will denote $\hat{q}$. With the calibration step done, CP sets can be generated for a test datapoint $x_{\text{test}}\sim\mathbb{P}$ as
$\mathcal{C}_{\hat{q}}(x_{\text{test}}):=\{y\in\mathcal{Y}\mid s(x_{\text{test}},y)<\hat{q}\}.$ (2)
For any classifier $f$ and any score function $s$, the CP sets $\mathcal{C}_{\hat{q}}$ will provide $1-\alpha$ coverage (Equation 1).
@@ 2.3 Non-conformal set predictors
$\mathcal{C}_{q_{k}}(x_{\text{test}}):=\{y\in\mathcal{Y}\mid f(x_{\text{test}})_{y}>q_{k}\}.$ (4)
@@ 2.4 Fairness of set predictors
$\delta_{t,a}:=\mathrm{acc}_{t}[x\in\mathcal{D}\mid g(x)=a]-\mathrm{acc}_{\mathrm{control}}[x\in\mathcal{D}\mid g(x)=a],$ (6)
@@
$\Delta_{t}:=\max_{a,b\in\mathcal{G}}(\delta_{t,a}-\delta_{t,b}),$ (7)
@@
$\Delta_{\text{Cov}}:=\max_{a,b\in\mathcal{G}}\left(\mathbb{P}[y\in\mathcal{C}(x)\mid g(x)=a]-\mathbb{P}[y\in\mathcal{C}(x)\mid g(x)=b]\right).$ (8)
@@ 4.1 Hypotheses
On its own this tradeoff is not directly connected to the substantive fairness of using prediction sets. However, it has been observed in previous studies that it is set size, not coverage, which has the greatest influence on human accuracy, and therefore on outcomes (Cresswell et al., 2024). Together, these facts imply that equalizing coverage will create more imbalanced set sizes which will propagate into more biased human performance. The proposed mechanism by which equalizing coverage with Mondrian (conditional) CP causes disparate impact is illustrated in Figure 2.
@@ 4.2 Method
Alongside empirical estimates of the standard errors, it is straightforward to derive odds ratios (ORs) from the GEE model. ORs allow us to estimate how much more likely a participant in treatment $t$ is to give the correct answer than if they were in the control, which is another way of expressing the expected accuracy improvement due to $t$ (c.f. Equation 6). For the sake of clarity, we give the formal description below. Let $p_{t,a}:=\hat{\mathbb{P}}(\mathrm{correct}=1\mid\mathrm{treat}=t,\mathrm{group}=a)$ be the probability of a correct answer under the GEE model, conditional on treatment and group. For treatment $t$, and for every protected group $a$, the OR of $t$ versus control is given by
$\mathbf{OR}_{t,a}:=\frac{p_{t,a}/(1-p_{t,a})}{p_{\mathrm{control},a}/(1-p_{\mathrm{control},a})}.$ (10)
@@
To assess the disparate impact caused by any treatment for Hypothesis 1, we can compare ORs between groups. We define the ratio of ORs (RORs) for treatment $t$ and groups $a$ and $b$ as $\mathbf{ROR}_{t,a,b}:=\mathbf{OR}_{t,a}/\mathbf{OR}_{t,b}$. When $\mathbf{ROR}_{t,a,b}\approx 1$, treatment $t$ provides the same advantage over the control for group $a$ as it does for group $b$ which is a fair outcome. Hence, a value of $\mathbf{ROR}_{t,a,b}$ greater than 1 for some pair $(a,b)$ indicates that $t$ causes disparate impact, and so we report the maximum ROR between pairs of groups as a descriptive statistic (c.f. Equation 7),
@@ 5.1 Tasks, Datasets, and Models
Additional details on tasks, datasets, models, and set prediction methods are given in Appendix A. Our code is available at github.com/layer6ai-labs/conformal-prediction-fairness.
@@ 5.2 Experiment Design
During each experiment, participants were first shown a consent form detailing how their data would be collected and used. Then, participants were introduced to the task and trained on 20 practice trials which were not used in our analysis. The testing phase proceeded with $M=50$ more trials, an example of which is shown in Figure 3 for FACET, while detailed task descriptions can be found in Appendix B. For each trial, participants were presented with a datapoint $x$ along with all $m$ class labels and were asked to classify $x$. For the avg-$k$, marginal, and conditional treatments participants were also shown a prediction set, accompanied by the expected coverage. There was no time limit for responses, however for FACET the image $x$ was only displayed for one second to increase task difficulty. After the class was selected for each trial, the correct answer was revealed.
@@ 6.1 Disparate Impact Measurement In Human Subject Experiments
Hypothesis 1 supposes that biases from a model can propagate to humans using prediction sets as an aid and result in disparate impact. Table 1 demonstrates that the models we used have some biases across groups, with non-zero $\Delta_{\text{Top-1}}$ values, which translated to non-zero $\Delta_{\text{Cov}}$ for the avg-$k$ and marginal sets. Conditional sets do equalize coverage ($\Delta_{\text{Cov}}\approx 0$), but at the cost of greater discrepancies in set size, $\Delta_{\text{Size}}$. From Table 2 we see that in most cases prediction sets were useful to participants versus the control ($\mathbf{OR}_{t,a}>1$), supporting previous findings (e.g. Cresswell et al. (2024)). However, it is clear that the beneficial effect is not experienced equally across groups. For FACET, the Younger group had the highest model accuracy (Figure 5), and experienced the most improvement across all treatments. The Older and Unknown groups having lower model accuracy and minority representation (see Table 6 in Appendix A) saw less improvement, and even some harm with conditional sets where human performance decreased ($\mathbf{OR}_{t,a}<1$). Hence, all three treatments caused disparate impact ($\mathbf{maxROR}_{t}>1)$. Similar results are seen for BiosBias. Even though the model itself was much less biased (Table 1) and both Female and Male groups benefited when sets were provided, the amount of benefit was not equally shared. Both FACET and BiosBias show that providing prediction sets can cause disparate impact, in support of Hypothesis 1. RAVDESS provides an example that disparate impact need not always occur. Again, both groups benefited from each treatment, but for the avg-$k$ and marginal treatments the increase actually was commensurate across groups. For these two treatments the prediction sets did not have equal coverage, but did have close to the same average set sizes across groups.
@@ 6.2 Insights
By contrast, we see strong correlation between set size differences and $\Delta_{t}$. The most fair outcomes occur when set sizes are similar across all groups, as for RAVDESS-Marginal, or FACET-Avg-K. The conditional treatment which increases set size differences in its effort to equalize coverage also increases disparate impact. Similar to set size, there is also clear correlation between singleton frequency and $\Delta_{t}$. Cresswell et al. (2024) observed that human accuracy on tasks was highest when the model expressed certainty through singleton sets. Hence, groups with more singletons could expect higher accuracy, so it is plausible that equalizing how often singletons occur would encourage fair outcomes, which is what we observe. Users should be aware that group-conditional modifications such as equalizing singleton frequency are susceptible to Yule effects (Ruggieri et al., 2024).
In summary, our results show that simply equalizing coverage, as is the aim of the conditional treatment, is not an effective way to promote fair outcomes. Instead, focusing on Equalized Set Size or Equalized Singleton Frequency could be more useful standards of fairness for set predictors.
@@ 7 Conclusions & Limitations
Ethics Statement Throughout our research, we took steps to ensure that research ethics were upheld. We acknowledge that ethical considerations are extremely important for any research involving human subjects, and we have ensured that our research meets the Code of Ethics for ICLR 2025. We also followed the ethical standards of our institution. Although it does not have an internal review board (IRB) process, the steps we took to meet its standards included: understanding the protocols and procedures for ethical research at our institution; curating datasets that only contained content suitable for showing to participants; performing the experiments on ourselves as authors; providing participants with a clear consent form that detailed the data that would be collected as well as how it would be used and how long it would be retained; paying all recruits a fair wage; and ensuring we were aware of potential biases in the data we collected by reviewing the demographic distribution of recruits (Table 8).
Reproducibility Statement We have made two specific efforts to promote the reproducibility of our results. First, we have released our codebase and the data we collected at github.com/layer6ai-labs/conformal-prediction-fairness, which will enable other researchers to perform the same set prediction methods, generate the same dataset splits, and perform the same statistical analysis. Second, we have extensively described our experimental setup in Appendix B including images of the screens that were shown to human participants throughout the experiment. The description is complete enough for a third party to faithfully recreate and reproduce our experiment.
@@ Appendix A Implementation Details
Our code for curating datasets and performing calibration and set prediction is available at github.com/layer6ai-labs/conformal-prediction-fairness. In this section we detail the steps that were taken.
@@ A.1 Dataset Pre-processing
For RAVDESS (Livingstone & Russo, 2018), a dataset of audio-visual recordings where actors express emotions, we worked with 1,440 audio-only samples exhibiting $m=8$ emotions: [Happy, Angry, Calm, Fearful, Neutral, Disgust, Sad and Surprised]. The dataset contains recordings from 24 actors, 12 female and 12 male, so we used emotion as the class label and gender as the protected group. Each actor recorded 60 snippets of approximately 4 seconds in length each. The 60 recordings are broken down as 2 re-recordings each of 2 intensities each of 2 statements of 8 emotions. However, the “Neutral” emotion used only a single intensity. We split the dataset into 240 for calibration-validation, 840 samples for calibration, and 360 for testing using stratified sampling over both emotion and gender. Given that each split was treated the same way they can be considered i.i.d. for conformal prediction. To process the audio samples, we used a wav2Vec2 model Baevski et al. (2020) available on Huggingface (Fadel, 2023). We resampled each file to 16kHz as the base model was pre-trained on that frequency. For audio with multiple channels, we changed it to mono by averaging channels. The model’s feature extractor then turned the raw audio signals into feature vectors ready for the model to use, which were subsequently padded to ensure consistent input length for batch processing by the model. RAVDESS is released under a Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International license.